OpenStudy (callisto):

Integration question #2

OpenStudy (callisto):

How should I start?

OpenStudy (anonymous):

Take out your pencil/pen. For a start.

OpenStudy (anonymous):

lol ishaan

OpenStudy (callisto):

@Ishaan94 Oh... Thank you....

OpenStudy (anonymous):

remember the COH- cos=0pp./hyp.

OpenStudy (anonymous):

\[I_2 =\int_0^{\pi}g(\pi-x) \ln(1 + e^{\cos(\pi - x)}) \]

OpenStudy (anonymous):

\[2I = \int_0^{\pi} -g(x)\ln(1 + e^{-\cos x}) + g(x)\ln(1+e^{\cos x})\] \[2I = \int _)^{\pi} g(x) \left(\ln \frac{1 + e^{\cos x}}{1 + \frac1{e^{\cos x}}}\right)\]

OpenStudy (anonymous):

\[I = \frac12 \int_0^{\pi}g(x)\cos x\]

OpenStudy (anonymous):

Watch this video, and only then you will become a JEDI Knight of Math. http://www.youtube.com/watch?v=vEanZ-Ex5Uw hehe ...kidding.

OpenStudy (anonymous):

If you don't understand which property @Ishaan94 used, Here it is \[\int\limits_{a}^{b}f(x)dx = \int\limits_{a}^{b}f(a+b-x)dx\]

OpenStudy (anonymous):

Yes. Yes.

OpenStudy (callisto):

Actually... In the question... It stated that using the substitution u=pi -x ...

OpenStudy (anonymous):

It's the same. You can try it on your own.

OpenStudy (callisto):

Oh!~ Finally I got it! Thanks !!!

OpenStudy (lgbasallote):

AWWW i missed it :/ is there a #3???

OpenStudy (callisto):

I understand Ishaan's answer. But I can't do it with u-sub:S Yea.. perhaps there is a #3

OpenStudy (anonymous):

ohhh NOOOOOOOOO, i lost my answer due to stupid google chrome crash. @callisto perform the substitution, I will guide you. I don't wanna type the answer again. Do the substitution and then tell me what you get.

OpenStudy (anonymous):

NICE ONE @Ishaan94 :))

OpenStudy (callisto):

\[I = \int_{0}^{\pi}g(x)\ln (1+e^{cosx})dx\]\[I = -\int_{0}^{\pi}g(\pi-x)\ln (1+e^{cos(\pi-x)})dx\] Let u = π - x du = -dx , when x=0, u=π ; when x=π, u=0\[I = \int_{\pi}^{0}g(u)\ln (1+e^{cos(u)})du\] Up to this step, any problem?

OpenStudy (anonymous):

oh thanks :-)

OpenStudy (anonymous):

-cosu not cosu

OpenStudy (anonymous):

wait...

OpenStudy (anonymous):

yes negative cosu

OpenStudy (callisto):

Why?

OpenStudy (anonymous):

cos(pi - x) =/= cos x but cos(pi - x) = -cos x

OpenStudy (callisto):

Hmm... if it is not u-sub I understand this concept. So, when we do u-sub, we need to change the sign also?

OpenStudy (callisto):

I meant if we need to consider the quad. it lies

OpenStudy (anonymous):

of course.

OpenStudy (anonymous):

waait... why did you change the limit?

OpenStudy (callisto):

u-sub....Shouldn't I change the limit?

OpenStudy (anonymous):

Ignore all my earlier advice and follow this. \[I=\int _0^{\pi} g(\pi - x) \ln (1 +e^{\cos(\pi-x)})dx\]After the U substitution\[I=-\int_0^{\pi}g(u)\ln(1+e^{\cos u})du\tag1\]From (1) \[I =-\int_0^{\pi} g(\pi -u)\ln (1+e^{\cos(\pi - u)})\implies I =\int_0^{\pi} g(u)\ln (1+e^{-\cos u})\tag2 \] Now add 1 and 2.

OpenStudy (callisto):

So, still we should do u-sub first. Then, replace x by u in the condition given ( or vice versa). Next, do something sensible, right?

OpenStudy (anonymous):

yeah you will have to replace the u with x.

OpenStudy (callisto):

Okay, that makes sense now. Thank you. And what's wrong with your previous solution?

OpenStudy (anonymous):

Nothing is wrong with my previous solution? (The Latex one)

OpenStudy (callisto):

Alright.... Now I understand the u-sub, not your previous one :S But thanks!